OpenStudy (anonymous):

Write the product as a sum. 11sin(x/2)cos(x/4)

OpenStudy (anonymous):

forget the 11 for moment and you use \[\sin(\alpha)\cos(\beta)=\frac{1}{2}(\sin(\alpha +\beta)+\sin(\alpha -\beta))\]

OpenStudy (anonymous):

in this case \[\alpha =\frac{x}{2},\beta=\frac{x}{4}\] so you can write this as \[\frac{11}{2}(\sin(\frac{3x}{4})+\sin(\frac{x}{4}))\]